OpenStudy (anonymous):

simplify the expression -l1-4l

OpenStudy (tkhunny):

Subtract First.

OpenStudy (anonymous):

be carful remember becuast it's n the bars all of it should be positive so it's actually -|1+4|

OpenStudy (anonymous):

it should simplify to -1+4

OpenStudy (anonymous):

3

OpenStudy (anonymous):

if im right may i have a medal and fan please :P

OpenStudy (tkhunny):

@topgunschield @AhhhhhplzzzzzHelp @me11 Ummm... NO! Absolute values do not alter subtraction. Absolute values also do not cancel the Distributive Property. -l1-4l As was stated in the beginning, SUBTRACT. 1 - 4 = -3, giving... -l-3l Absolute values make things positive. |-3| = 3 -(3) And we're done. -3 Let's be a little more careful, shall we?

OpenStudy (anonymous):

it is -3 positive 3 is wrong.

OpenStudy (anonymous):

but you dont subtract first you make it positive first and then subtract. You cant have negatives in absolute value bars

OpenStudy (tkhunny):

Here's the plan, please learn rather than argue. It is correct as I have demonstrated it. Please back up and work 50 problems in Absolute Value. You will see that you have not stated it correctly. -|1-4| = -|-3| = -(3) = -3 There is nothing wrong with negatives inside absolute values. When you EVALUATE the absolute value, it had better be positive. BEFORE you evaluate it, there is nothing to do. Subtraction is not altered by an external absolute value.

OpenStudy (anonymous):

There is when you have to make the negatives positive. Goodbye I have others to help